### v1

# Vamos: Versatile Action Models for Video Understanding

Abstract
What makes good video representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as discrete action labels, or free-form video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularities. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embeddings, action labels, and free-form descriptions extracted from videos as its input. We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, Next-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We perform extensive ablation study and qualitative analysis to support our observations, and achieve state-of-the-art performance on three benchmarks. Code and models will be released at https://brown-palm.github.io/Vamos/.
1 Introduction
Figure 1: Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations, discrete action labels, and free-form text descriptions for downstream applications.
A picture held us captive … for it lay in our language and language seemed to repeat it to us inexorably.
Philosophical Investigations
Building a generative model for everyday human activities has long been a dream for researchers working on video understanding. Central to this problem are capturing the interactions between humans and the environment [67, 12], modeling the temporal dynamics of activities [65, 15], and encoding the hierarchical structures among atomic actions [59, 23], activities [6, 7], and events [37, 35]. Once constructed, the generative model of actions can be applied to a wide range of tasks, including activity and event recognition [56], future behavior prediction [49], goal and intent inference [57], and temporal reasoning [71].
Despite its desirable properties, generative modeling of actions from video observations remains challenging, hindered by two open research questions: First, what makes good video representations? Earlier attempts often relied on manually defining the actions and the objects being interacted with [23, 1, 59]. They require task-specific prior knowledge, and cannot generalize to the “open vocabulary” scenarios in the wild. Alternative approaches aim to model the temporal dynamics of human pose [51, 45, 36] or latent representations encoded by deep neural networks [65, 32], which are either too fine-grained, or not directly intepretable. Second, what makes a good model of human actions? While earlier approaches attempted to apply rule-based generative action grammars [30, 25, 54, 56], they may not be able to capture the diverse, even peculiar ways of how events would unfold over time. More recent approaches adopt a data-driven framework and directly learn autoregressive models [40] on visual tokens [62, 74], where the visual domain is often specialized (e.g. cooking, or robotics).
We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [41, 48] on sampled video frames, serve as useful video representations for action modeling from videos? And if so, can we then leverage a pre-trained large language model (LLM) [64] as the generative model of actions, represented as free-form text? Despite the apparent crudeness, we argue that it is inspired and supported by recent advances in LLM research, where LLMs have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [22, 52], and performing reasoning [4, 13, 70, 14], all of which are indispensable for action modeling.
To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings, discrete action labels, and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [64]. The visual embeddings are linearly projected into the same language space following standard practice [41, 53]. As illustrated in Figure 1, the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type. Vamos can directly leverage an LLM’s next token prediction capability for action anticipation [21]. Similar to how an LLM can be tasked to perform reasoning, we ask Vamos to perform video question answering [71], by prepending the question to the video representation.
We perform extensive evaluations on four benchmarks, including the Ego4D [21] long-term action anticipation, NeXT-QA [71] and IntentQA [43] video question answering, and EgoSchema [50] zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the representation based on discrete action labels outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [68] by 50%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings only result in marginal performance gains.
Another benefit of text-based video representation is interpretability. We exploit this property in two scenarios: First, we perform causal intervention and correct the suboptimal video descriptions, and we observe that the answers are indeed corrected accordingly. Second, we propose a lightweight token selector module to sample as few as 20 tokens from video descriptions that are more than 600 tokens long. We observe that not only the video question answering performance remains comeptitive, outperforming the visual embedding baseline, but the selected tokens are indeed highly related to answering the questions, confirming that the Vamos framework behaves as intended.
To summarize, we make the following contributions:
  1. 1.
    We propose to revisit video representation and action modeling in the LLM era, and to explore the effectiveness of free-form text descriptions.
  2. 2.
    We propose Vamos, a versatile framework that allows us to incorporate and compare the effectiveness of different video representations, both visual and textual.
  3. 3.
    We demonstrate the effectiveness of free-form text as a representation that is performant, interpretable, and can be intervened, even when aggressively compressed to 6%. Vamos achieves state-of-the-art performance on Ego4D LTA, NeXT-QA, IntentQA, and EgoSchema.
2 Related Work
Figure 2: Overview of Vamos and token selector. (a) Vamos flexibly unifies distributed visual features and textual video representations including discrete action labels and free-form video captions. By performing sequence modeling, Vamos can be applied to multiple video understanding tasks. (b) The lightweight token selector as an add-on module of Vamos. It pick one single token from the video sequence based on the task sequence. By applying the token selector to $k$ segments of textual video representation, $k$ tokens are finally selected as the compact input to the LLM for down-stream tasks.
Vision-Language Foundation Models. Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension. Recent advancements in vision-language models have been propelled by various pretraining techniques leveraging large-scale multimodal datasets. Models such as CLIP [58], ALIGN [33], CLOOB [19] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs. Another line of vision-language models showcased by VisualBERT [44], UNITER [11], OSCAR [46], FLAVA [61], uses a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection. For videos, VIOLET [17] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [62, 17, 81, 82]. The objectives can be combined [78] and the encoders for different modalities can be shared [66]. Compared to existing vision-language models, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
Visually-augmented LLMs. Apart from joint visual-language pre-training, existing large language models (LLMs) can also be augmented to incorporate visual inputs. For example, VisualGPT [9] and Flamingo [2] directly fuse visual information into the layers of a language model decoder using a cross-attention mechanism instead of using images as additional prefixes to the language model. Other approaches, such as instructional tuning [48], or linearly projecting the visual embeddings into the input space of LLMs [41, 53], have also been explored. Vamos largely follows this linear projection approach to incorporate visual embedding inputs, with the goal to understand if and how they are complementary to text-based video representations.
Additionally, tool-using large language models have been recently proposed to invoke and incorporate the use of task-specific modules [60, 27], where visual perceptions consist a substantial subset of the tools. Notably, VisualProgram [24] and ViperGPT [63] propose to apply LLMs to generate symbolic programs based on pre-selected computer vision modules for visual question answering. Closest to our work is Socratic Models [83], where the authors propose to use natural language as the common interface to connect foundation models with different input and output modalities. Vamos differs from these approaches as it aims to understand the power and limitations of text-based representation for a wide range of video understanding tasks, through quantitative experiments and qualitative analysis.
3 Method
We now describe how the text-based representation is constructed and incorporated into versatile action models.
3.1 Text-based Video Representation
Compared with a single image, the video data often contains more complex and dynamic information including context and interactions. Prior works [39, 80, 69] have demonstrated the effectiveness of condensing images into conceptual or natural language bottleneck with good interpretability, while it is still unclear if long form videos with richer dynamics can also be condensed into such explainable textual representations. To translate videos into language bottleneck, we consider two methods: discrete action labels and free-form text descriptions.
We propose to use an extremely compact representation by condensing videos into sequences of discrete action labels, offering a more compact representation. In practice, this is achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
Video text descriptions use free-form natural language to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video. It transforms complex and dynamic visual information into more comprehensible and interpretable textual descriptions. We employ off-the-shelf image captioning models such as BLIP-2 [42] that process sampled frames to produce frame-level captions, which are subsequently concatenated to form a comprehensive video-level caption.
3.2 Versatile Action Models
Benefiting from the encoded extensive prior knowledge, large language models (LLMs) have demonstrated strong capability for temporal and causal reasoning, which are crucial for video understanding. We introduce Vamos, a simple yet effective framework to utilize LLMs to unify video dynamic modeling tasks, including comprehending historical content (video question answering, VQA) and future prediction (long-term action anticipation, LTA). As shown in Figure 2 (a), given a video $V$ and a pretrained LLM $\pi$, the input sequence $x_{t}=[x_{tvr},x_{task}]$ consists of the textual video representations $x_{tvr}$ of $V$ and other task specific langauge inputs $x_{task}$ (e.g., instructions, questions, targets). The frozen word embedding layer $\mathcal{F}_{emb}$ first generate the corresponding text tokens $z_{t}=\mathcal{F}_{emb}(x_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $x_{t}$, $D$ is the feature dimension.
To address any residual information not entirely captured by the text-based representations, we incorporate a learnable linear projection layer $\mathcal{F}_{proj}$ to align visual features with the language space. Specifically, the frozen vision backbone $\mathcal{E}$ takes in $N_{v}$ frames $[v_{1},...v_{N_{v}}]$ sampled from $V$ to generate the visual features. These visual features are then input into the projection layer $\mathcal{F}_{proj}$ to produce vision tokens $z_{v}=\mathcal{F}_{proj}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$. $z_{v}$ and $z_{t}$ are then concatenated and fed into the LLM $\pi$. During training, the LLM $\pi$ is optimized with the standard language modeling loss using LoRA [26].
Vamos is accommodating multiple video understanding tasks that can be formatted as sequence modeling tasks with varying task-specific inputs $x_{task}$ and modeling objectives. Specifically, for the VQA task, $x_{task}$ is composed of instructions, questions, and answers, with the answer being the training objective. During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA, depending on the dataset. For the LTA task, $x_{task}$ is composed of instructions and future actions, where the training objective is the future action sequence. During inference, the trained LLM generates future action sequences based on history actions.
3.3 Lightweight Token Selector
We also investigate the potential of compressing long input sequences and extracting keys element from the generated textual representations to probe the crucial information for downstream video understanding tasks. We design a lightweight token selector as a add-on module (shown in Figure 2 (b)) for Vamos. The token selector takes in a sequence of textual video tokens to be selected, and a task-specific token sequence, functioning as queries. It picks a single token from the video tokens that is most relevant for downstream video understanding. To select a condensed token sequence for a video, the original textual video token sequence with $L$ tokens is first uniformly divided into $k\ll L$ segments $\{X^{(1)}$,…, $X^{(k)}\}$ containing $n$ tokens. Each segment $X^{(i)}=\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ is fed into the token selector as candidate tokens, from which one token $x^{(i)}$ is selected based on the task texts (question for VQA).
Within the token selector, $\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ are first projected to a lower dimension, then input into a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$. A linear layer then takes these encodings and generates the logits $g^{(i)}\in\mathbb{R}^{n}$ for final selection. During training, to maintain model differentiability, we apply Gumbel-Softmax [31] to the logits $g^{(i)}$ to pick the final selected token $x^{(i)}$ for each segment $X^{(i)}$. In this way, $k$ tokens $\{x^{(1)},...,x^{(k)}\}$ are sampled as a condensed representation of the original text sequence with much smaller length and input to downstream LLMs for video understanding.
4 Experiments
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Input Ego4D-LTA normal-$\downarrow$
Verb Noun Action
vision 0.653 0.673 0.884
action 0.661 0.651 0.878
caption 0.676 0.675 0.890
vis+act 0.643 0.650 0.868
Input NeXT-QA normal-$\uparrow$ IntentQA normal-$\uparrow$
Cau. Tem. Des. All CW CH TP&TN All
vision 61.5 58.2 66.3 61.2 55.2 62.4 46.9 54.4
action 55.1 53.4 53.2 54.3 57.5 65.2 47.4 56.3
caption 70.9 69.3 77.7 71.5 69.7 68.8 61.7 67.6
vis+cap 70.9 67.2 77.7 70.8 69.0 69.9 63.2 67.7
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Table 2: Vamos with various video representations on VQA task. Results are reported on NeXT-QA and IntentQA test set, the metric is accuracy. On both dataset, caption-based representations achieve competitive performance.
In this section, we conduct experiments on two tasks and four datasets with both quantitative and qualitative analysis.
4.1 Task and Datasets
We focus on two major video understanding tasks: long-term action anticipation and video question answering. Long-term action anticipation. The long-term action anticipation (LTA) task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions. In LTA, a long video $V$ is first split into a number of annotated video segments. Given video observation before segment $i$, our task is to predict the future actions in sequences of verb-noun pairs of the next $Z$ segments allowing $K$ candidate sequences. The correctness of the predicted sequence is measured with edit distance. We evaluation on:
Ego4D [21] is comprised of 3,670 hours of egocentric videos in hundreds of scenarios of daily life activity. The Ego4D LTA v2 benchmark we focus on includes a total duration of around 243 hours of videos annotated into 3472 clips with 117 verbs and 521 nouns. We follow the official dataset splits and adopt the official parameters of the evaluation metric, with $Z=20$ and $K=5$.
Video question answering. Given a set of videos $V$, and a corresponding set of language-based questions $Q_{v}$ and their candidate answers $A_{q}$. The goal of video question answering (VQA) task is to predict the correct answer $A$ for each video-question pair. The performance is measured by accuracy. For VQA, we evaluate on three datasets:
EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding. In contrast to previous video question answering benchmarks where the input videos are generally less than one-minute long, Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos. It contains 5031 videos and each video has a multiple choice question. It requires zero-shot evaluation and releases answers to 500 questions for development.
NeXT-QA [71] is a popular multiple choice video question answering benchmark that test video understanding in terms of describing and reasoning the temporal actions. It contains 5,440 video clips and 47,692 questions, grouped into causal (48%), temporal (29%) and descriptive (23%).
IntentQA [43] is another multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning. The authors select the videos related to causal and temporal questions from NeXT-QA and constructed their own questions and answers to focus on testifying models’s performance on reasoning questions.
4.2 Implementation
Generating Action Labels. To generate action labels for videos, we use a recognition model pretrained on Ego4D LTA. It is a 3-block 6-head transformer-encoder that takes 4 CLIP features and outputs two logits for verb and noun respectively. It predicts actions in the action space pre-defined by Ego4D LTA. For each video, our action recognition model samples 4 frames for each 8s segment splits uniformly and output a verb and noun pair for the segment.
Generating Video Captions. We zero-shot generate task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema, LLaVA-1.5 for NeXT-QA. For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, then caption each frame and concatenate these frame-level captions to represent the video. For NeXT-QA, IntentQA, and EgoSchema we sample 6, 6, and 12 frames respectively.
Vamos for Temporal Modeling. For full-shot VQA and LTA, We use Llama2-7B [64], an open-sourced LLM as the temporal model for video understanding. During training, we use low-rank adaption [26] (LoRA) adapter to perform parameter-efficient fine-tuning on the training set. For vision input, we use the frozen CLIP [58] ViT-L/14 to extract image features. For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4 and Llama2-chat-13B.
4.3 LTA: Action Modeling with Vamos
We first apply Vamos on the long-term action anticipation task, which requires direct modeling of video temporal dynamics by predicting future actions tokens based on video observation. We finetune Vamos on Ego4D LTA dataset with three types of representations: the recognized action labels, free-formed captions, and the continuous vision features. As detailed in the first three rows in Table 2, we observe that the action-based representation typically outperforms the vision-based input. Notably, the action representation, despite its compact nature, surpasses captions, which provide more extensive details and descriptions of the observed events. This suggests that a discrete and structured sequence of actions serves as an effective representation for modeling video temporal dynamics.
4.4 LLM as Long-term Video Temporal Reasoner
Our experiments on LTA task reveal that compact action labels are effective for video temporal modeling, but is action representation alone sufficient for comprehensive video understanding? We then delve into a more challenging task: long-form video question answering, which demands an extensive understanding and reasoning of the video content. We conduct zero-shot experiment on EgoSchema, a recent challenging video question answering benchmark where existing state-of-the-art visual-language models struggle.
Model Input Type Full Set Acc.
InternVideo [68] frame 32.1%
GPT-4 action 38.12%
GPT-4 caption 48.26%
GPT-4* gt-narration 81.80%
Table 3: zero-shot VQA on Egoschema. * denotes subset results.
We employ OpenAI GPT-4 as the zero-shot video reasoner, and the same models used for LTA to zero-shot extract action labels and captions (BLIP-2) as well. Results are listed in Table 3. We observe the LLM utilizing both zero-shot generated actions and captions largely outperform current best-performed visual-language model (InternVideo [68]). Contrasting with the action prediction task, captions with abundant details and semantic information serve as a much stronger representation for the challenging VQA task. Furthermore, we use the video narrations of 50 examples provided by the authors as “oracle” captions and evaluate on this subset. Remarkably, the LLM achieves an impressive accuracy of 81.8% with the “oracle” captions. Although not directly comparable with the full-set performance, it reflects the vast potential and generalizability of caption-based textual video representations and the LLMs’ capability for reasoning over broad time spans.
4.5 What Makes Good Video Representation?
In addition to text features, Vamos can also integrate vision-language features. We then investigate whether different modalities encode complementary information on Ego4D, IntentQA, and Next-QA. We train the Vamos model to incorporate the visual embeddings (the linear layer in Figure 2), and use LoRA to update the weights of Llama-2.
From Ego4D LTA results shown in Table 2, we observe incorporating vision input with the compact action sequence further enhances model performance especially on verb and action prediction. For the VQA benchmarks NeXT-QA and IntentQA, as illustrated in Table 2, the caption-based textual video representations significantly outperform purely visual representations in both content description and causal reasoning. This aligns with our prior observations from EgoSchema that captions can serve as a strong representation for question answering and reasoning. Contrasting with the LTA task, adding visual features to captions only marginally influence the performance on the two VQA datasets. This suggests that the task-agnostic captions already contain sufficient information and cues to address the questions in these benchmarks.
4.6 How many tokens are needed for reasoning?
While we show zero-shot captions generated by open-box models are effective representations of video and offer a competitive interface to interact with LLMs, we further ask the following questions: How many tokens are really necessary to represent videos? What specific information from lengthy captions is actually being utilized by LLMs? We then implant a lightweight token selector to Vamos to automatically compress the captions into much more compact token sequences.
We conduct experiment on the NeXT-QA benchmark. In detail, after the word embedding layer in Vamos, the long caption-based textual video token sequence is uniformly partitioned into 40 segments. The token selector are then applied on each segment to select one token. In this way, we condense a long sequence (644 on average) into 40 tokens. These 40 tokens are then fed into LLMs as the new compressed representation of original video for down-stream video question answering task. During training, the token selector is optimized together with the Vamos.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. We show the results in Table 4. After condensing the prompts from an average of 644 tokens to 40 tokens, Vamos experiences an expected performance drop but still achieves a competitive 65.3% accuracy, significantly outperforming the vision-only baseline utilizing double input frames by 4.1%. When further compressed to 20 tokens, the performance drops to 59.5%, slightly worse than the vision baseline. This shows the sparsity of the free-form natural language captions for video reasoning and indicates a promising potential to compress the textual representation for better practical application. Besides, instead of using hard selection (choosing one token with highest score) during inference, we also perform soft-selection (using weighted pooling by applying Softmax instead of hard-max to the logits). Interestingly, the soft-selection achieves similar performance with hard-selection even with access to the whole sequence, indicating there is few missing information left in the unselected tokens.
 Input selected # Cau. Tem. Des. All
Frame all 61.5 58.2 66.3 61.2
Caption all 70.9 69.3 77.7 71.5
Caption 20 / 644 59.5 56.8 64.6 59.5
Caption 40 / 644 65.4 62.4 70.5 65.3
Caption* 40 / 644 65.6 62.4 70.4 65.4
Table 4: Condensing captions with token selector on NeXT-QA. Average caption length is 644. * denotes soft-selection results.
Figure 3: Illustration of Vamos predictions and manual intervention.
Figure 4: Illustration of predictions with token selector and manual intervention. The selector picks tokens highly related to the question and intervention corrects wrong predictions without training.
4.7 Qualitative Analysis and Intervention
Compared with continuous visual features, one major advantage of the textual video representation is it’s excellent interpretability that making understanding the relationship between input and prediction direct and simple. We then make qualitative analysis of some positive and negative results from Vamos and the token selector.
Visualization of results. In Figure 3 we demonstrate two correctly-answered examples and one incorrect of Vamos from IntentQA. From the two positive examples we can see that the generated captions well describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”). However, a suboptimal caption will also cause the Vamos’ failure of reasoning. In the third negative example, the captions successfully describe the baby’s action of eating but fail to acknowledge the presence and potential actions of a woman in the corner of the scene. This omission leads to an incorrect prediction regarding the woman’s reaction.
Visualization of Token Selector. By condensing the captions consists of thousands of tokens into 40 with a token selector, it not only significantly accelerates the training and inference speed, but also provides a better view for us to understand what’s the key element for Vamos to understand and reason about the video. In Figure 4, we demonstrate two positive examples and one negative example of Vamos with a token selector from NeXT-QA. After the token selector, captions in natural language are broken into separated words that highly descriptive and related to the question. The two positive examples demonstrate Vamos’ strong ability to reason by the selected tokens providing direct or indirect clues to the question. In the challenging negative example, even highly related tokens such as “race” and “track” are selected, Vamos still tends to relate “tyres” with the more common word “car”, indicating the potential bias encoded by the LLMs from pre-training.
Test Time Intervention. In addition to good interpretability, the text-based white-box representation also provides the capability of post-hoc test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining the reasoning model. In Figure 3 and Figure 4, we show examples of test-time intervention of the negative examples for both original Vamos and Vamos with selector. By providing more accurate and related captions or tokens, Vamos is able to correct the wrong answers.
4.8 Design Choices and Ablation Study
Caption models. We study the impact of caption models on the reasoning performance on NeXT-QA. We compare two captioning models: BLIP-2 and LLaVA-1.5-7b. We observe that captions generated by BLIP-2 are generally concise spanning less than 20 tokens long while their counterparts from LLaVA-1.5 are more detailed and descriptive with about 100 tokens on average. Results shown in Table 5 shows that captions from LLaVA-1.5 achieve better performance. Interestingly, the two captions achieve similar performance on descriptive questions but LLaVA-1.5 captions perform significantly better on casual and temporal reasoning. This indicates that reasoning from videos generally benefit more from extra details compared with describing and summarizing.
Number of Captions. We study the impact of sampled frame numbers for captioning on NeXT-QA. As shown in Table 5, we found that with more captions of frames spanning across the video the model will perform better. This result is align with our expectation as we bring information with more granularity from the videos.
Comparison of LLMs on zero-shot VQA Besides, we also compare various LLMs’ zero-shot performance based on the same caption representations on EgoSchema. As shown in Table 6, GPT-4 achieves huge improvements comparing with GPT-3.5-turbo and Llama2-Chat-7B, demonstrating the potentiality and advantage of textual representation to naturally incorporate the improvements of LLMs.
 Caption Frame # Cau. Tem. Des. All
LLaVA-1.5 1 66.4 63.0 72.0 66.3
LLaVA-1.5 3 68.5 65.2 76.4 68.8
LLaVA-1.5 6 70.9 69.3 77.7 71.5
BLIP-2 6 67.9 64.6 77.1 68.4
Table 5: Ablation on captioning model and frame numbers.
4.9 Comparison with State-of-the-art
Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 678 and 9. On EgoSchema zero-shot VQA, our approach based on GPT-4 outperforms the best vision-language model by 50.3%. On NeXT-QA, Vamos achieves the best performance among baselines with similiar model size. It is slightly worse than SeViLA, which trains on additional dataset for temporal localization. On IntentQA, Vamos outperforms all baselines, with a 17.4% accuracy improvement. On Ego4d LTA, Vamos outperforms AntGPT without relying on domain-specific video encoders [47].
Model Input Type Frames
VIOLET [18] frame 19.9%
FrozenBiLM [75] frame 26.9%
mPLUG-Owl [77] frame 31.1%
InternVideo [68] frame 32.1%
Llama2-13B caption 36.73%
GPT-3.5 caption 41.24%
GPT-4 caption 48.26%
Table 6: Egoschema VQA zero-shot Performance.
 Model Cau. Tem. Des. All
ATP [5] 53.1% 50.2% 66.8% 54.3%
MIST [20] 54.6% 56.6% 66.9% 57.2%
HiTeA [76] 62.4% 58.3% 75.6% 63.1%
ATM [10] 55.3% 55.6% 65.3% 57.0%
Intern Video [68] 62.5% 58.5% 75.8% 63.2%
BLIP-2 [42] 70.1% 65.2% 80.1% 70.1%
LLaMA-VQA-7B [38] 72.7% 69.2% 75.8% 72.0%
Vamos (ours) 72.6% 69.6% 78.0% 72.5%
SeViLA[79]* 74.2% 69.4% 81.3% 73.8%
Table 7: Comparison on NeXT-QA benchmark. * denotes model using additional dataset during training.
 Model CW CH TP&TN ALL
HGA [34] 44.88% 50.97% 39.62% 44.61%
HQGA [72] 48.24% 54.32% 41.71% 47.66%
VGT [73] 51.44% 55.99% 47.62% 51.27%
BlindGPT [55] 52.16% 61.28% 43.43% 51.55%
CaVIR [43] 58.4% 65.46% 50.48% 57.64%
Vamos (ours) 69.52% 70.19% 64.95% 68.51%
Table 8: Comparison with SOTA on IntentQA
 Model verb noun action
Slowfast [16] 0.717 0.736 0.925
VideoLLM [8] 0.721 0.725 0.921
PaMsEgoAI [29] 0.684 0.679 0.893
Palm [28] 0.696 0.651 0.886
AntGPT [84] 0.650 0.650 0.877
Vamos (ours) 0.643 0.650 0.868
Table 9: Comparison with SOTA on Ego4D v2 test set
5 Conclusion
We study the form of representation for long videos and proposed versatile action models (Vamos) for video understanding. Through extensive experiments on long-term action anticipation and video question answer benchmarks, we have demonstrated that surprisingly, direct application of free-form text descriptions generated by vision-language models serve as strong video representation for all benchmarks we consider, and the use of large language models enable us to perform zero-shot reasoning, and correct the answers via causal intervention. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, close to SOTA performance on NeXT-QA, and outperforms the best vision-language model by over 50% on EgoSchema. Although our results show the promise of free-form text representation, we believe visual information is still essential for complex video understanding and reasoning, and that further explorations on modeling and benchmarking are needed.
Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their generous feedback on this work. Our research was conducted using computational resources at Brown CCV.

### v2

@@
 Abstract
-What makes good video representations for video understanding, such as anticipating future activities, or answering video-conditioned questions?
-While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as discrete action labels, or free-form video captions, which are interpretable and can be directly consumed by large language models (LLMs).
-Intuitively, different video understanding tasks may require representations that are complementary and at different granularities.
-To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embeddings, action labels, and free-form descriptions extracted from videos as its input.
+What makes good representations for video understanding, such as anticipating future activities, or answering video-conditioned questions?
+While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as general-purpose video captions, which are interpretable and can be directly consumed by large language models (LLMs).
+Intuitively, different video understanding tasks may require representations that are complementary and at different granularity.
+To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embedding and free-form text descriptions as its input.
+To interpret the important text evidence for question answering, we generalize the concept bottleneck model to work with tokens and nonlinear models, which uses hard attention to select a small subset of tokens from the free-form text as inputs to the LLM reasoner.
 We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, NeXT-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning.
 Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era.
-We perform extensive ablation study and qualitative analysis to support our observations, and achieve state-of-the-art performance on three benchmarks.
-Code and models will be released at https://brown-palm.github.io/Vamos/.
+We also demonstrate that our token bottleneck model is able to select relevant evidence from free-form text, support test-time intervention, and achieves nearly 5 times inference speedup while keeping a competitive question answering performance.
+Code and models are publicly released at https://brown-palm.github.io/Vamos/.
 1 Introduction
-Figure 1: Visual observations with vastly diverse appearances may be described with the same captions.
-Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering.
-We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations, discrete action labels, and free-form text descriptions for downstream applications.
-A picture held us captive … for it lay in our language and language seemed to repeat it to us inexorably.
-Philosophical Investigations
 Building a generative model for everyday human activities has long been a dream for researchers working on video understanding.
@@
 cooking, or robotics).
+Figure 1: (a) Visual observations with vastly diverse appearances may be described with the same captions.
+Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering.
+We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations and free-form text descriptions for downstream applications.
+(b) Overview of the Vamos framework.
+It flexibly unifies distributed visual features and textual video representations such as free-form video captions.
+By performing sequence modeling, Vamos can be applied to multiple video understanding tasks.
 We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [40, 48] on sampled video frames, serve as useful video representations for action modeling from videos?
 And if so, can we then leverage a pre-trained large language model (LLM) [67] as the generative model of actions, represented as free-form text?
-Despite the apparent crudeness, we argue that it is inspired and supported by recent advances in LLM research, where LLMs have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [22, 52], and performing reasoning [4, 13, 70, 14], all of which are indispensable for action modeling.
-To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings, discrete action labels, and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [64].
+If both answers are yes, visual reasoning can then enjoy recent advances in LLM research, where they have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [19, 53], and performing reasoning [4, 12, 74, 13], all of which are indispensable for action modeling.
+Since we assume the text descriptions are general purpose, they can be extracted once and reused for different downstream tasks, similar to pre-computed visual embeddings [59].
+To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [67].
 The visual embeddings are linearly projected into the same language space following standard practice [40, 54].
-As illustrated in Figure 1, the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type.
+As illustrated in Figure 1(a), the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type.
 Vamos can directly leverage an LLM’s next token prediction capability for action anticipation [18].
-Similar to how an LLM can be tasked to perform reasoning, we ask Vamos to perform video question answering [71], by prepending the question to the video representation.
-We perform extensive evaluations on four benchmarks, including the Ego4D [21] long-term action anticipation, NeXT-QA [71] and IntentQA [43] video question answering, and EgoSchema [50] zero-shot long-form video question answering.
-We observe that for the direct application of Vamos in the action anticipation task, the representation based on discrete action labels outperforms its counterpart based on visual embeddings.
+We also ask Vamos to perform video question answering [75], by appending the question to the video representation as inputs to the LLM reasoner.
+One inherent benefit of text-based video representation is its interpretability.
+Inspired by interpretable object classifiers such as concept bottleneck models (CBM) [37], we aim to understand which words serve as important evidence for question answering.
+However, CBM requires a pre-defined list of discrete visual concepts, and requires a linear classifier to achieve interpretability.
+We generalize this framework and learns hard attention to select a small subset of text inputs to feed to the LLM reasoner, where the text inputs are tokenized text as opposed to pre-defined concepts.
+We call our generalized formulation token bottleneck models (TBM).
+TBM naturally supports the incorporation of multimodal information, and allows users to perform causal intervention.
+We perform extensive evaluations on four benchmarks, including the Ego4D dataset [18] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, and EgoSchema [50] for zero-shot long-form video question answering.
+We observe that for the direct application of Vamos in the action anticipation task, the text-based representation outperforms its counterpart based on visual embeddings.
 We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 50%.
-We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings only result in marginal performance gains.
-Another benefit of text-based video representation is interpretability.
-We exploit this property in two scenarios: First, we perform causal intervention and correct the suboptimal video descriptions, and we observe that the answers are indeed corrected accordingly.
-Second, we propose a lightweight token selector module to sample as few as 20 tokens from video descriptions that are more than 600 tokens long.
-We observe that not only the video question answering performance remains comeptitive, outperforming the visual embedding baseline, but the selected tokens are indeed highly related to answering the questions, confirming that the Vamos framework behaves as intended.
-To summarize, we make the following contributions:
-  1. 1.
-    We propose to revisit video representation and action modeling in the LLM era, and to explore the effectiveness of free-form text descriptions.
-  2. 2.
-    We propose Vamos, a versatile framework that allows us to incorporate and compare the effectiveness of different video representations, both visual and textual.
-  3. 3.
-    We demonstrate the effectiveness of free-form text as a representation that is performant, interpretable, and can be intervened, even when aggressively compressed to 6%.
-Vamos achieves state-of-the-art performance on Ego4D LTA, NeXT-QA, IntentQA, and EgoSchema.
+We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings surprisingly results in marginal or no performance gains.
+Finally, we demonstrate that the token bottleneck model is able to select semantically relevant evidence for question answering, and achieves 5x speedup at inference time while maintaining the question answering accuracy.
 2 Related Work
-Figure 2: Overview of Vamos and token selector.
-(a) Vamos flexibly unifies distributed visual features and textual video representations including discrete action labels and free-form video captions.
-By performing sequence modeling, Vamos can be applied to multiple video understanding tasks.
-(b) The lightweight token selector as an add-on module of Vamos.
-It pick one single token from the video sequence based on the task sequence.
-By applying the token selector to $k$ segments of textual video representation, $k$ tokens are finally selected as the compact input to the LLM for down-stream tasks.
 Vision-Language Foundation Models.
 Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension.
-Recent advancements in vision-language models have been propelled by various pretraining techniques leveraging large-scale multimodal datasets.
 Models such as CLIP [59], ALIGN [30], CLOOB [17] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs.
-Another line of vision-language models showcased by VisualBERT [44], UNITER [11], OSCAR [46], FLAVA [61], uses a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection.
+Another line of vision-language models [43, 10, 45, 63] use a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection.
 For videos, VIOLET [16] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling.
@@
 The objectives can be combined [82] and the encoders for different modalities can be shared [69].
-Compared to existing vision-language models, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
+Compared to existing VLMs, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
 Visually-augmented LLMs.
@@
 For example, VisualGPT [9] and Flamingo [2] directly fuse visual information into the layers of a language model decoder using a cross-attention mechanism instead of using images as additional prefixes to the language model.
-Other approaches, such as instructional tuning [48], or linearly projecting the visual embeddings into the input space of LLMs [41, 53], have also been explored.
-Vamos largely follows this linear projection approach to incorporate visual embedding inputs, with the goal to understand if and how they are complementary to text-based video representations.
+Other approaches, such as instructional tuning [48], prompting large language models for knowledge retrieval [62], or linearly projecting the visual embeddings into the input space of LLMs [54, 46], have also been explored.
+Vamos largely follows this approach to incorporate visual embeddings, with the goal to understand if and how they are complementary to text-based video representations.
 Additionally, tool-using large language models have been recently proposed to invoke and incorporate the use of task-specific modules [61, 24], where visual perceptions consist a substantial subset of the tools.
 Notably, VisualProgram [21] and ViperGPT [65] propose to apply LLMs to generate symbolic programs based on pre-selected computer vision modules for visual question answering.
+VidIL [72] leverages expert knowledge to design object and action concepts for few-shot captioning and video question answering.
 Closest to our work is Socratic Models [87], where the authors propose to use natural language as the common interface to connect foundation models with different input and output modalities.
-Vamos differs from these approaches as it aims to understand the power and limitations of text-based representation for a wide range of video understanding tasks, through quantitative experiments and qualitative analysis.
+Finally, several concurrent works [88, 42, 51] share similar motivations and methods to Vamos in utilizing text-based representations and modules for video understanding, without the incorporation of visual embeddings.
 3 Method
@@
 3.1 Text-based Video Representation
-Compared with a single image, the video data often contains more complex and dynamic information including context and interactions.
-Prior works [39, 80, 69] have demonstrated the effectiveness of condensing images into conceptual or natural language bottleneck with good interpretability, while it is still unclear if long form videos with richer dynamics can also be condensed into such explainable textual representations.
-To translate videos into language bottleneck, we consider two methods: discrete action labels and free-form text descriptions.
-We propose to use an extremely compact representation by condensing videos into sequences of discrete action labels, offering a more compact representation.
-In practice, this is achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
-Video text descriptions use free-form natural language to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video.
-It transforms complex and dynamic visual information into more comprehensible and interpretable textual descriptions.
-We employ off-the-shelf image captioning models such as BLIP-2 [42] that process sampled frames to produce frame-level captions, which are subsequently concatenated to form a comprehensive video-level caption.
+A video often contains complex and dynamic information including context and interactions.
+While prior works [37, 84, 73] have demonstrated the effectiveness of condensing images into text-based representations such as visual concepts, it remains unclear if videos can also be condensed into text-based representations.
+To answer this research question, we consider text descriptions that are task-agnostic, and can potentially be applied in diverse video understanding tasks.
+Concretely, we rely on general-purpose captioning models to generate free-form text descriptions to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video.
+We employ off-the-shelf image captioning models such as BLIP-2 [41] that generate image-level captions from the sampled video frames.
+These captions are subsequently concatenated to form a comprehensive video-level caption.
+For certain tasks, prior knowledge might be helpful to guide the model learning.
+For example, when the goal is to model the long-term temporal dynamics of verbs and nouns for the long-term action anticipation task, it would be beneficial to trim the inputs to only contain discrete action labels.
+In practice, this can be achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
 3.2 Versatile Action Models
-Benefiting from the encoded extensive prior knowledge, large language models (LLMs) have demonstrated strong capability for temporal and causal reasoning, which are crucial for video understanding.
+Large language models (LLMs) have demonstrated strong capability for temporal reasoning [90] and even some potential for causal reasoning [35], both of which are crucial for video understanding.
 We introduce Vamos, a simple yet effective framework to utilize LLMs to unify video dynamic modeling tasks, including comprehending historical content (video question answering, VQA) and future prediction (long-term action anticipation, LTA).
-As shown in Figure 2 (a), given a video $V$ and a pretrained LLM $\pi$, the input sequence $x_{t}=[x_{tvr},x_{task}]$ consists of the textual video representations $x_{tvr}$ of $V$ and other task specific langauge inputs $x_{task}$ (e.g., instructions, questions, targets).
-The frozen word embedding layer $\mathcal{F}_{emb}$ first generate the corresponding text tokens $z_{t}=\mathcal{F}_{emb}(x_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $x_{t}$, $D$ is the feature dimension.
-To address any residual information not entirely captured by the text-based representations, we incorporate a learnable linear projection layer $\mathcal{F}_{proj}$ to align visual features with the language space.
+As shown in Figure 1 (b), given a video $V$ and a pretrained LLM $\pi$, the input sequence $\mathbf{x}_{t}=[\mathbf{x}_{\text{tvr}},\mathbf{x}_{\text{task}}]$ consists of the textual video representations $\mathbf{x}_{\text{tvr}}$ of $V$ and other task specific language inputs $\mathbf{x}_{\text{task}}$ (e.g., instructions, questions, targets).
+The frozen word embedding layer $\mathcal{F}_{\text{emb}}$ first generate the corresponding text tokens $\mathbf{z}_{t}=\mathcal{F}_{\text{emb}}(\mathbf{x}_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $\mathbf{x}_{t}$, $D$ is the feature dimension.
+Vamos incorporates the residual information not entirely captured by $\mathbf{x}_{\text{tvr}}$ via representations encoded directly from the visual modality, such as CLIP visual embedding.
+We adopt a learnable linear projection layer $\mathcal{F}_{\text{proj}}$ to align visual features with the language space.
 Specifically, the frozen vision backbone $\mathcal{E}$ takes in $N_{v}$ frames $[v_{1},...v_{N_{v}}]$ sampled from $V$ to generate the visual features.
-These visual features are then input into the projection layer $\mathcal{F}_{proj}$ to produce vision tokens $z_{v}=\mathcal{F}_{proj}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$.
-$z_{v}$ and $z_{t}$ are then concatenated and fed into the LLM $\pi$.
-During training, the LLM $\pi$ is optimized with the standard language modeling loss using LoRA [26].
-Vamos is accommodating multiple video understanding tasks that can be formatted as sequence modeling tasks with varying task-specific inputs $x_{task}$ and modeling objectives.
-Specifically, for the VQA task, $x_{task}$ is composed of instructions, questions, and answers, with the answer being the training objective.
-During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA, depending on the dataset.
-For the LTA task, $x_{task}$ is composed of instructions and future actions, where the training objective is the future action sequence.
-During inference, the trained LLM generates future action sequences based on history actions.
-3.3 Lightweight Token Selector
-We also investigate the potential of compressing long input sequences and extracting keys element from the generated textual representations to probe the crucial information for downstream video understanding tasks.
-We design a lightweight token selector as a add-on module (shown in Figure 2 (b)) for Vamos.
-The token selector takes in a sequence of textual video tokens to be selected, and a task-specific token sequence, functioning as queries.
-It picks a single token from the video tokens that is most relevant for downstream video understanding.
-To select a condensed token sequence for a video, the original textual video token sequence with $L$ tokens is first uniformly divided into $k\ll L$ segments $\{X^{(1)}$,…, $X^{(k)}\}$ containing $n$ tokens.
-Each segment $X^{(i)}=\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ is fed into the token selector as candidate tokens, from which one token $x^{(i)}$ is selected based on the task texts (question for VQA).
-Within the token selector, $\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ are first projected to a lower dimension, then input into a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$.
-A linear layer then takes these encodings and generates the logits $g^{(i)}\in\mathbb{R}^{n}$ for final selection.
-During training, to maintain model differentiability, we apply Gumbel-Softmax [31] to the logits $g^{(i)}$ to pick the final selected token $x^{(i)}$ for each segment $X^{(i)}$.
-In this way, $k$ tokens $\{x^{(1)},...,x^{(k)}\}$ are sampled as a condensed representation of the original text sequence with much smaller length and input to downstream LLMs for video understanding.
+These visual features are then fed into the projection layer $\mathcal{F}_{\text{proj}}$ to produce visual tokens $\mathbf{z}_{v}=\mathcal{F}_{\text{proj}}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$.
+To combine information from the visual and textual representations, we adopt the early fusion strategy and concatenate $\mathbf{z}_{v}$ and $\mathbf{z}_{t}$ as the inputs to the LLM $\pi$.
+When labeled training data is available for task-specific fine-tuning, we update the weights of the LLM $\pi$ either with LoRA [23] or LLaMA-Adapter [89].
+Vamos can accommodate diverse video understanding tasks by formulating each task as sequence completion given an appropriate task description $\mathbf{x}_{\text{task}}$, the LLM $\pi$ can then be optimized with the standard language modeling objectives.
+Specifically, for the VQA task, $\mathbf{x}_{\text{task}}$ is composed of instructions, questions, and answers, with the answer being the training objective.
+During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA.
+For the LTA task, $\mathbf{x}_{\text{task}}$ is composed of instructions and future actions, where the training objective is the future action sequence.
+During inference, the fine-tuned LLM is tasked to generate sequences of future actions based on the history actions.
+Discussion: By design, Vamos naturally incorporates temporal information as the captions are timestamped.
+Since the text-based representation is general-purpose, the framework is efficient since once extracted, the text-based representation can be reused for different questions, just like the CLIP embeddings commonly used by VLMs.
+Vamos is a two-stage framework with decoupled “perception” and “reasoning” modules, in addition to the benefits on interpretability (of the intermediate representation) and efficiency (reuse of the intermediate features).
+Another conceptual benefit is its generalizability: While the visual distributions of different datasets may differ, the reasoning module may be shared.
+We show this is indeed the case for EgoSchema [50], when the end-to-end vision-language models performance significantly worse than Vamos.
+3.3 Token Bottleneck Models
+Figure 2: An illustration of the token bottleneck model (TBM).
+We are inspired by the concept bottleneck models (CBM) [37], which achieve interpretable object classification by inspecting the weights of the learned linear classifier (left).
+Unlike CBM, Vamos does not require pre-defining a list of concepts.
+It directly works with tokenized text inputs.
+To provide input tokens to the reasoning model (an LLM), we leverage hard attention to generate binary rather than continuous weights (middle).
+The token bottleneck can be interpreted directly.
+It can also be intervened with human inputs (right), or augmented with residual visual information.
+We aim to understand how the text-based representations are utilized by the LLM reasoners.
+As an LLM operates like a black box, we aim to enhance the interpretability of the overall framework by first understanding how it selects evidence to solve the downstream tasks.
+As illustrated in Figure 2, we are inspired by the success of interpretable object classifiers, such as the concept bottleneck model (CBM) [37].
+CBM relies on pre-defining a list of concepts and building (often supervised) concept detectors for each of them.
+As our model’s inputs are free-form text, we propose to directly work with word tokens as opposed to pre-defined concepts.
+In addition, CBM relies on linear classifiers to achieve model interpretability.
+Each weight in a learned classifier indicates the importance of the corresponding concept for making the prediction.
+We hypothesize that linear classifiers are not sufficiently expressive when solving tasks that require a stronger “reasoner”.
+To strike a balance between model interpretability and expressiveness, we generalize the CBM framework to learn binary attention on the input tokens as opposed to continuous weights used by the CBM linear classifier.
+As illustrated in Figure 2, The binary weights indicate which tokens are to be selected and fed to the more expressive LLM for solving the target tasks.
+We name this generalized framework as token bottleneck models (TBMs).
+To implement TBM, we design a lightweight token selector as an add-on module for Vamos (Figure 2 middle).
+The token selector takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs.
+It learns to pick a single token among the candidate tokens $\mathbf{z}_{\text{tvr}}$ by optimizing the objective for a given downstream task.
+To select a sequence of tokens, we assume the important information is even distributed across the input sequence, and uniformly divide the input sequence into $k$ segments $\{\mathbf{z}_{\text{tvr}}^{(1)}$,…, $\mathbf{z}_{\text{tvr}}^{(k)}\}$, each of which contains $n$ tokens.
+Each segment $\mathbf{z}_{\text{tvr}}^{(i)}=\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ is fed into the token selector, from which one token $z^{(i)}$ is selected for the task $\mathbf{z}_{\text{task}}$.
+Within the token selector, $\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ are first projected to a lower dimension, and then provided as inputs to a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$.
+A linear layer then takes these encodings and generates the logits $\mathbf{g}^{(i)}\in\mathbb{R}^{n}$ for final selection.
+During training, we apply Gumbel-Softmax [28] on the logits $\mathbf{g}^{(i)}$ to pick the token $z^{(i)}$ for each segment $\mathbf{z}_{\text{tvr}}^{(i)}$ while ensuring the module is differentiable.
+In this way, $k$ tokens are selected as the condensed representation of the original tokenized input sequence $\mathbf{z}_{\text{tvr}}$.
+The token selector in TBM allows us to inspect the important evidence selected for the downstream tasks, and to intervene the wrongly recognized or selected tokens with the correct ones with human in the loop (Figure 2 right).
+Practically, the token selector also can also speed up the inference time due to its own light-weight implementation, and that only a much smaller subset of the tokens (e.g.
+6%) are processed by the computationally heavy LLM.
 4 Experiments
-Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
-Input Ego4D-LTA normal-$\downarrow$
-Verb Noun Action
-vision 0.653 0.673 0.884
-action 0.661 0.651 0.878
-caption 0.676 0.675 0.890
-vis+act 0.643 0.650 0.868
-Input NeXT-QA normal-$\uparrow$ IntentQA normal-$\uparrow$
-Cau. Tem. Des. All CW CH TP&TN All
-vision 61.5 58.2 66.3 61.2 55.2 62.4 46.9 54.4
-action 55.1 53.4 53.2 54.3 57.5 65.2 47.4 56.3
-caption 70.9 69.3 77.7 71.5 69.7 68.8 61.7 67.6
-vis+cap 70.9 67.2 77.7 70.8 69.0 69.9 63.2 67.7
-Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
-Table 2: Vamos with various video representations on VQA task.
-Results are reported on NeXT-QA and IntentQA test set, the metric is accuracy.
-On both dataset, caption-based representations achieve competitive performance.
 In this section, we conduct experiments on two tasks and four datasets with both quantitative and qualitative analysis.
 4.1 Task and Datasets
-We focus on two major video understanding tasks: long-term action anticipation and video question answering.
 Long-term action anticipation.
-The long-term action anticipation (LTA) task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions.
+The LTA task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions.
 In LTA, a long video $V$ is first split into a number of annotated video segments.
@@
 EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding.
-In contrast to previous video question answering benchmarks where the input videos are generally less than one-minute long, Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos.
+Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos.
 It contains 5031 videos and each video has a multiple choice question.
 It requires zero-shot evaluation and releases answers to 500 questions for development.
-NeXT-QA [71] is a popular multiple choice video question answering benchmark that test video understanding in terms of describing and reasoning the temporal actions.
+NeXT-QA [75] is a popular multiple choice video question answering benchmark that tests video understanding in terms of describing and reasoning the temporal actions.
 It contains 5,440 video clips and 47,692 questions, grouped into causal (48%), temporal (29%) and descriptive (23%).
-IntentQA [43] is another multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning.
+IntentQA [39] is a multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning.
 The authors select the videos related to causal and temporal questions from NeXT-QA and constructed their own questions and answers to focus on testifying models’s performance on reasoning questions.
@@
 Generating Video Captions.
-We zero-shot generate task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema, LLaVA-1.5 for NeXT-QA.
+We generate zero-shot, task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema.
+We use LLaVA-1.5 to generate captions for NeXT-QA due to its better performance (see Table 4.5 for an ablation study).
 For Ego4D LTA, we sample the center frame for each video segment to generate its caption.
 The captions for the 8 observed segments are then concatenated as the video representations.
-For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, then caption each frame and concatenate these frame-level captions to represent the video.
-For NeXT-QA, IntentQA, and EgoSchema we sample 6, 6, and 12 frames respectively.
+For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, and then generate and concatenate the frame-level captions.
+We sample 6, 6, and 12 frames for NeXT-QA, IntentQA, and EgoSchema, respectively.
 Vamos for Temporal Modeling.
-For full-shot VQA and LTA, We use Llama2-7B [64], an open-sourced LLM as the temporal model for video understanding.
-During training, we use low-rank adaption [26] (LoRA) adapter to perform parameter-efficient fine-tuning on the training set.
+For full-shot VQA and LTA, We use LLaMA-7B and LLaMA2-7B [67] respectively as the temporal model for video understanding.
+During training, we use LLaMA Adapter [89] or low-rank adaption [23] (LoRA) to perform parameter-efficient fine-tuning on the training set.
 For vision input, we use the frozen CLIP [59] ViT-L/14 to extract image features.
-For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4 and Llama2-chat-13B.
-4.3 LTA: Action Modeling with Vamos
+For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4, GPT-4o and LLaMA2-chat-13B.
+Table 1: Vamos with various video representations.
+Results are reported on Ego4D LTA test set, the metric is edit distance.
+Results are reported on NeXT-QA validation set, and IntentQA test set, the evaluation metric is accuracy.
+On all datasets, text-based representations achieve competitive performance.
+Input Ego4D-LTA $\downarrow$ NeXT-QA $\uparrow$ IntentQA $\uparrow$
+Verb Noun Action Cau. Tem. Des. All CW CH TP&TN All
+vision 0.653 0.673 0.884 69.6 67.2 74.7 69.6 68.9 71.6 58.0 66.7
+text 0.661 0.651 0.878 75.5 71.3 81.1 75.0 74.0 78.6 67.5 73.2
+vis+text 0.643 0.650 0.868 74.5 71.0 81.7 74.5 73.5 76.6 64.3 71.7
+4.3 LLM as Long-term Video Temporal Reasoner
 We first apply Vamos on the long-term action anticipation task, which requires direct modeling of video temporal dynamics by predicting future actions tokens based on video observation.
-We finetune Vamos on Ego4D LTA dataset with three types of representations: the recognized action labels, free-formed captions, and the continuous vision features.
-As detailed in the first three rows in Table 2, we observe that the action-based representation typically outperforms the vision-based input.
-Notably, the action representation, despite its compact nature, surpasses captions, which provide more extensive details and descriptions of the observed events.
-This suggests that a discrete and structured sequence of actions serves as an effective representation for modeling video temporal dynamics.
-4.4 LLM as Long-term Video Temporal Reasoner
-Our experiments on LTA task reveal that compact action labels are effective for video temporal modeling, but is action representation alone sufficient for comprehensive video understanding?
-We then delve into a more challenging task: long-form video question answering, which demands an extensive understanding and reasoning of the video content.
-We conduct zero-shot experiment on EgoSchema, a recent challenging video question answering benchmark where existing state-of-the-art visual-language models struggle.
-Model Input Type Full Set Acc.
+We fine-tune Vamos on Ego4D LTA dataset with continuous visual embeddings or text-based representation.
+We observe that action-based representation slightly outperforms the free-form captions due to the nature of the task.
+As shown in the first three columns of Table 4.2, we observe that the text-based representation outperforms the vision-based input, and combining the two further improves the performance.
+Table 2: zero-shot VQA on Egoschema. *: 500 question subset.
+Model Input Type   Acc.
 InternVideo [71] frame 32.1%
-GPT-4 action 38.12%
-GPT-4 caption 48.26%
+GPT-4 text 48.26%
 GPT-4* gt-narration 81.80%
-Table 3: zero-shot VQA on Egoschema. * denotes subset results.
-We employ OpenAI GPT-4 as the zero-shot video reasoner, and the same models used for LTA to zero-shot extract action labels and captions (BLIP-2) as well.
-Results are listed in Table 3.
-We observe the LLM utilizing both zero-shot generated actions and captions largely outperform current best-performed visual-language model (InternVideo [68]).
-Contrasting with the action prediction task, captions with abundant details and semantic information serve as a much stronger representation for the challenging VQA task.
-Furthermore, we use the video narrations of 50 examples provided by the authors as “oracle” captions and evaluate on this subset.
-Remarkably, the LLM achieves an impressive accuracy of 81.8% with the “oracle” captions.
-Although not directly comparable with the full-set performance, it reflects the vast potential and generalizability of caption-based textual video representations and the LLMs’ capability for reasoning over broad time spans.
+Table 3: Ablation on the number of frames on EgoSchema with GPT-3.5 turbo.
+# Frames   Full Set Acc.
+1 37.83%
+4 38.36%
+12 41.24%
+We now consider the more challenging task, long-form video question answering, by conducting zero-shot experiment on the recently collected EgoSchema benchmark.
+We employ OpenAI GPT-4 as the video “reasoner”.
+We observe that free-form captions extracted by BLIP-2 consistently outperform the action-based representation, presumably because richer information is retained in the captions.
+We report the free-form caption-based performance from now on, unless otherwise mentioned.
+From Table 4.3, we observe that Vamos with text-based video representation largely outperforms the state-of-the-art InternVideo model [71], which is jointly trained on vision and language inputs.
+We attribute the performance gain due to the decoupling of perception and reasoning, the latter of which we hypothesize is easier to generalize even at zero-shot, thanks to the LLM pre-training.
+To better understand the performance upper-bound of the text-based representation, we use 500 ground-truth video narrations provided by the authors of EgoSchema to perform evaluation on this subset.
+Remarkably, the LLM achieves an impressive accuracy of 81.8% with the oracle captions.
+Although not directly comparable with the full-set performance, the result confirms the potential of the Vamos framework for reasoning over broad time spans in the long-form video question answering task, and that better empirical performance may be achieved by improving the captioning models.
 4.4 What Makes Good Video Representation?
@@
 We then investigate whether different modalities encode complementary information on Ego4D, IntentQA, and NeXT-QA.
-We train the Vamos model to incorporate the visual embeddings (the linear layer in Figure 2), and use LoRA to update the weights of Llama-2.
-From Ego4D LTA results shown in Table 2, we observe incorporating vision input with the compact action sequence further enhances model performance especially on verb and action prediction.
-For the VQA benchmarks NeXT-QA and IntentQA, as illustrated in Table 2, the caption-based textual video representations significantly outperform purely visual representations in both content description and causal reasoning.
-This aligns with our prior observations from EgoSchema that captions can serve as a strong representation for question answering and reasoning.
-Contrasting with the LTA task, adding visual features to captions only marginally influence the performance on the two VQA datasets.
-This suggests that the task-agnostic captions already contain sufficient information and cues to address the questions in these benchmarks.
-4.6 How many tokens are needed for reasoning?
-While we show zero-shot captions generated by open-box models are effective representations of video and offer a competitive interface to interact with LLMs, we further ask the following questions: How many tokens are really necessary to represent videos?
-What specific information from lengthy captions is actually being utilized by LLMs?
-We then implant a lightweight token selector to Vamos to automatically compress the captions into much more compact token sequences.
-We conduct experiment on the NeXT-QA benchmark.
-In detail, after the word embedding layer in Vamos, the long caption-based textual video token sequence is uniformly partitioned into 40 segments.
-The token selector are then applied on each segment to select one token.
-In this way, we condense a long sequence (644 on average) into 40 tokens.
-These 40 tokens are then fed into LLMs as the new compressed representation of original video for down-stream video question answering task.
-During training, the token selector is optimized together with the Vamos.
-For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA.
+We perform parameter efficient fine-tuning to update the weights of the LLM, whose gradients are used to jointly train the linear projection layer to incorporate the visual embeddings.
+We observe that naively fine-tuning with text and visual inputs lead to model overfitting, and hence perform modality dropout (i.e.
+randomly discard the entire sequence of visual embeddings) when fine-tuning the vis+text models.
+While visual and text inputs are complementary in the Ego4D LTA task, we observe in Table 4.2 that the caption-based representation significantly outperform the vision-only baseline in the NeXT-QA and IntentQA benchmarks for video question answering, and that adding visual features only marginally affects the performance.
+This suggests that pre-trained visual embeddings such as CLIP may not encode residual information useful for the video QA task (e.g.
+fine-grained visual and motion features).
+Extracting visual representations that would complement the information encoded by the task-agnostic captions is thus an important future work.
+Figure 3: Visualizations of example Vamos predictions and manual intervention.
+4.5 Token Bottleneck Models
+Although the text-based video representation is directly interpretable, it remains unclear which tokens are selected as evidence for making predictions.
+The token bottleneck models (TBMs) enable us to reveal the selected evidence and also improve the model’s inference speed.
+VQA Performance.
+We first tokenize and embed the input sequence $\mathbf{x}_{t}$ to obtain $\mathbf{z}_{t}$, which is uniformly partitioned into $k=$ 20 or 40 segments.
+TBM is then applied on each segment to select one token.
+In this way, we condense a long sequence (644 tokens on average for NeXT-QA) into $k$ tokens.
+TBM is jointly optimized with Vamos during training.
+For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA-1.5.
 For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input.
-We show the results in Table 4.
-After condensing the prompts from an average of 644 tokens to 40 tokens, Vamos experiences an expected performance drop but still achieves a competitive 65.3% accuracy, significantly outperforming the vision-only baseline utilizing double input frames by 4.1%.
-When further compressed to 20 tokens, the performance drops to 59.5%, slightly worse than the vision baseline.
-This shows the sparsity of the free-form natural language captions for video reasoning and indicates a promising potential to compress the textual representation for better practical application.
-Besides, instead of using hard selection (choosing one token with highest score) during inference, we also perform soft-selection (using weighted pooling by applying Softmax instead of hard-max to the logits).
-Interestingly, the soft-selection achieves similar performance with hard-selection even with access to the whole sequence, indicating there is few missing information left in the unselected tokens.
- Input selected # Cau. Tem. Des. All
-Frame all 61.5 58.2 66.3 61.2
-Caption all 70.9 69.3 77.7 71.5
-Caption 20 / 644 59.5 56.8 64.6 59.5
-Caption 40 / 644 65.4 62.4 70.5 65.3
-Caption* 40 / 644 65.6 62.4 70.4 65.4
-Table 4: Condensing captions with token selector on NeXT-QA. Average caption length is 644. * denotes soft-selection results.
-Figure 3: Illustration of Vamos predictions and manual intervention.
-Figure 4: Illustration of predictions with token selector and manual intervention.
-The selector picks tokens highly related to the question and intervention corrects wrong predictions without training.
-4.7 Qualitative Analysis and Intervention
-Compared with continuous visual features, one major advantage of the textual video representation is it’s excellent interpretability that making understanding the relationship between input and prediction direct and simple.
-We then make qualitative analysis of some positive and negative results from Vamos and the token selector.
-Visualization of results.
-In Figure 3 we demonstrate two correctly-answered examples and one incorrect of Vamos from IntentQA.
-From the two positive examples we can see that the generated captions well describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”).
-However, a suboptimal caption will also cause the Vamos’ failure of reasoning.
-In the third negative example, the captions successfully describe the baby’s action of eating but fail to acknowledge the presence and potential actions of a woman in the corner of the scene.
+Table 4.5 shows that the TBM leads to an expected performance drop due to discarding over 90% of the input tokens, and increasing $k$ improves the TBM performance.
+When $k=40$, Vamos achieves a competitive 65.3% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
+Figure 4: Illustration of the token bottleneck model and the impact of intervention on model’s predictions.
+The selected tokens are highly related to the task (examples $a$ and $b$) and can be manually intervened to correct the model’s prediction (example $c$).
+Table 4: Condensing captions with token bottleneck models on NeXT-QA.
+Input Selected tokens ($k$) Cau. Tem. Des. All
+vision all 69.6 67.2 74.7 69.6
+text all 75.5 71.3 81.1 75.0
+text 20 / 644 59.5 56.8 64.6 59.5
+text 40 / 644 65.4 62.4 70.5 65.3
+Visualization of Vamos predictions.
+Figure 3 provides two positive and one negative examples selected from IntentQA.
+The predictions are made by Vamos without TBM.
+From the two positive examples we can see that the generated captions manage to describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”).
+However, a sub-optimal caption will also cause the Vamos’s failure of reasoning.
+In the third example, we observe that the captions successfully describe the baby’s action of eating but fail to describe the presence and potential actions of a woman in the corner of the scene.
 This omission leads to an incorrect prediction regarding the woman’s reaction.
-Visualization of Token Selector.
-By condensing the captions consists of thousands of tokens into 40 with a token selector, it not only significantly accelerates the training and inference speed, but also provides a better view for us to understand what’s the key element for Vamos to understand and reason about the video.
-In Figure 4, we demonstrate two positive examples and one negative example of Vamos with a token selector from NeXT-QA.
-After the token selector, captions in natural language are broken into separated words that highly descriptive and related to the question.
-The two positive examples demonstrate Vamos’ strong ability to reason by the selected tokens providing direct or indirect clues to the question.
-In the challenging negative example, even highly related tokens such as “race” and “track” are selected, Vamos still tends to relate “tyres” with the more common word “car”, indicating the potential bias encoded by the LLMs from pre-training.
+Visualization of the Selected Tokens.
+TBM not only helps accelerate Vamos’ inference speed, but also allows us to interpret which pieces of evidence are used by the model to make predictions.
+In Figure 4, we show two positive examples and one negative example of Vamos with TBM from NeXT-QA.
+The two positive examples demonstrate Vamos’ ability to select tokens that serve as evidence for question answering.
+The negative example is challenging: even highly related tokens such as “race” and “track” are selected by Vamos, it still fails to reason the purposes of the tyres in the video.
 Test Time Intervention.
-In addition to good interpretability, the text-based white-box representation also provides the capability of post-hoc test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining the reasoning model.
-In Figure 3 and Figure 4, we show examples of test-time intervention of the negative examples for both original Vamos and Vamos with selector.
-By providing more accurate and related captions or tokens, Vamos is able to correct the wrong answers.
+Text-based video representation is not only directly interpretable as used by Vamos and TBM, it also allows the users to conduct post-hoc, test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining.
+In Figure 3 and Figure 4, we show examples of test-time intervention performed on the negative examples for Vamos and Vamos with TBM, respectively.
+By providing more accurate and related captions or tokens, Vamos is able to correct its wrong predictions.
+Table 5: Ablation on captioning model and frame numbers on NeXT-QA.
+ Caption # Frame Cau. Tem. Des. All
+LLaVA-13B 1 69.9 67.2 74.8 69.8
+LLaVA-13B 3 73.0 70.4 79.4 73.1
+LLaVA-13B 6 75.5 71.3 81.1 75.0
+LLaVA-7B 6 75.2 71.9 81.6 75.1
+BLIP-2 6 72.7 68.9 78.8 72.4
+Table 6: Comparison of different LLaMA models on NeXT-QA.
+Model Cau. Tem. Des. All
+LLaMA1-7B 75.5 71.3 81.1 75.0
+LLaMA2-7B 74.8 72.3 81.6 75.0
+LLaMA3-8B 77.2 75.3 81.7 77.3
 4.6 Design Choices and Ablation Study
 Caption Models.
-We study the impact of caption models on the reasoning performance on NeXT-QA.
-We compare two captioning models: BLIP-2 and LLaVA-1.5-7b.
-We observe that captions generated by BLIP-2 are generally concise spanning less than 20 tokens long while their counterparts from LLaVA-1.5 are more detailed and descriptive with about 100 tokens on average.
-Results shown in Table 5 shows that captions from LLaVA-1.5 achieve better performance.
-Interestingly, the two captions achieve similar performance on descriptive questions but LLaVA-1.5 captions perform significantly better on casual and temporal reasoning.
-This indicates that reasoning from videos generally benefit more from extra details compared with describing and summarizing.
-Number of Captions.
+We study the impact of caption models on the video QA performance on NeXT-QA.
+We compare two captioning models: BLIP-2 and LLaVA-1.5.
+We observe that captions generated by BLIP-2 are generally more concise (less than 20 tokens), while captions generated by LLaVA-1.5 are more detailed (around 100 tokens on average).
+Results shown in Table 4.5 shows that captions from LLaVA-1.5 achieve better performance.
+We also investigate the influence of caption model size by comparing LLaVA-1.5 7B and 13B versions.
+Interestingly, scaling LLaVA from 7B to 13B does not lead to improvement.
+Number of Frames.
 We study the impact of sampled frame numbers for captioning on NeXT-QA.
-As shown in Table 5, we found that with more captions of frames spanning across the video the model will perform better.
-This result is align with our expectation as we bring information with more granularity from the videos.
-Comparison of LLMs on zero-shot VQA Besides, we also compare various LLMs’ zero-shot performance based on the same caption representations on EgoSchema.
-As shown in Table 6, GPT-4 achieves huge improvements comparing with GPT-3.5-turbo and Llama2-Chat-7B, demonstrating the potentiality and advantage of textual representation to naturally incorporate the improvements of LLMs.
- Caption Frame # Cau. Tem. Des. All
-LLaVA-1.5 1 66.4 63.0 72.0 66.3
-LLaVA-1.5 3 68.5 65.2 76.4 68.8
-LLaVA-1.5 6 70.9 69.3 77.7 71.5
-BLIP-2 6 67.9 64.6 77.1 68.4
-Table 5: Ablation on captioning model and frame numbers.
+As shown in Table 4.5, we found that using more captioned frames leads to better performance.
+A similar trend can be observed on EgoSchema in Table 4.3.
+However, we observe diminishing return when 12 frames are used, and hence not worth the speed and accuracy trade-off.
+Impact of LLMs on VideoQA.
+We study the impact of different LLaMA versions on NeXT-QA.
+As shown in Table 4.5, Vamos directly benefits from a more advanced LLM, which is a desirable property for practitioners.
+Impact of LLMs on Long-form VideoQA.
+We study the impact of LLMs on the EgoSchema benchmark.
+As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines.
+This again demonstrates that Vamos can directly benefit from advances in the large language models, making its decoupled perception and reasoning design desirable for practitioners.
 4.7 Comparison with State-of-the-art
-Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 678 and 9.
-On EgoSchema zero-shot VQA, our approach based on GPT-4 outperforms the best vision-language model by 50.3%.
-On NeXT-QA, Vamos achieves the best performance among baselines with similiar model size.
-It is slightly worse than SeViLA, which trains on additional dataset for temporal localization.
-On IntentQA, Vamos outperforms all baselines, with a 17.4% accuracy improvement.
-On Ego4d LTA, Vamos outperforms AntGPT without relying on domain-specific video encoders [47].
-Model Input Type Frames
-VIOLET [18] frame 19.9%
-FrozenBiLM [75] frame 26.9%
-mPLUG-Owl [77] frame 31.1%
-InternVideo [68] frame 32.1%
-Llama2-13B caption 36.73%
-GPT-3.5 caption 41.24%
-GPT-4 caption 48.26%
-Table 6: Egoschema VQA zero-shot Performance.
+Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 4.74.74.7, and 4.7.
+We train Vamos with LLaMA2-7B on Ego4D LTA and LLaMA3-8B on NeXT-QA and IntentQA.
+On EgoSchema zero-shot VQA, our approach based on GPT-4o outperforms the best vision-language model by 66.8% on the full set.
+On NeXT-QA, Vamos achieves the best performance and significantly outperforms LLaMA-VQA [36] and SeViLA[83], even though the former uses a LLM with 33B parameters and the latter is trained on additional dataset with temporal localization supervision.
+On IntentQA, Vamos also outperforms all baselines, with a 28.7% accuracy improvement compared to the best performing prior method.
+On Ego4D LTA, Vamos outperforms previous works, without relying on domain-specific video encoders [47].
+Table 7: Comparison on NeXT-QA benchmark. * with additional supervision.
  Model Cau. Tem. Des. All
 ATP [5] 53.1% 50.2% 66.8% 54.3%
-MIST [20] 54.6% 56.6% 66.9% 57.2%
 HiTeA [80] 62.4% 58.3% 75.6% 63.1%
-ATM [10] 55.3% 55.6% 65.3% 57.0%
 Intern Video [71] 62.5% 58.5% 75.8% 63.2%
 BLIP-2 [41] 70.1% 65.2% 80.1% 70.1%
+SeViLA*[83] 74.2% 69.4% 81.3% 73.8%
 LLaMA-VQA-7B [36] 72.7% 69.2% 75.8% 72.0%
-Vamos (ours) 72.6% 69.6% 78.0% 72.5%
-SeViLA[79]* 74.2% 69.4% 81.3% 73.8%
-Table 7: Comparison on NeXT-QA benchmark. * denotes model using additional dataset during training.
+LLaMA-VQA-33B [36] 76.2% 72.6% 78.8% 75.5%
+Vamos (ours) 77.2% 75.3% 81.7% 77.3%
+Table 8: Comparison with SOTA on IntentQA
  Model CW CH TP&TN ALL
@@
 CaVIR [39] 58.4% 65.46% 50.48% 57.64%
-Vamos (ours) 69.52% 70.19% 64.95% 68.51%
-Table 8: Comparison with SOTA on IntentQA
+Vamos (ours) 75.14% 77.44% 69.58% 74.16%
+Table 9: Comparison with SOTA on Ego4D LTA v2 test set.
  Model verb noun action
@@
 Vamos (ours) 0.643 0.650 0.868
-Table 9: Comparison with SOTA on Ego4D v2 test set
+Table 10: Egoschema VQA zero-shot performance on full set and subset.
+Model Input Type Full Subset
+FrozenBiLM [79] frame 26.9% -
+mPLUG-Owl [81] frame 31.1% -
+InternVideo [71] frame 32.1% -
+Vamos (LLaMA2-13B) caption 36.73% 38.20%
+Vamos (GPT-3.5) caption 41.24% 47.60%
+Vamos (GPT-4) caption 48.26% 51.20%
+Vamos (GPT-4o) caption 53.55% 57.20%
 5 Conclusion
-We study the form of representation for long videos and proposed versatile action models (Vamos) for video understanding.
-Through extensive experiments on long-term action anticipation and video question answer benchmarks, we have demonstrated that surprisingly, direct application of free-form text descriptions generated by vision-language models serve as strong video representation for all benchmarks we consider, and the use of large language models enable us to perform zero-shot reasoning, and correct the answers via causal intervention.
-Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, close to SOTA performance on NeXT-QA, and outperforms the best vision-language model by over 50% on EgoSchema.
-Although our results show the promise of free-form text representation, we believe visual information is still essential for complex video understanding and reasoning, and that further explorations on modeling and benchmarking are needed.
+We study different forms of video representations and propose versatile action models (Vamos) as a unified framework to utilize visual- and text-based representations for video understanding.
+We conduct extensive experiments on long-term action anticipation and video question answer benchmarks.
+Surprisingly, we observe that direct applications of free-form, general-purpose text-based video representations, such as captions, serve as strong video representation for all benchmarks we consider.
+Vamos utilizes large language models to perform zero-shot reasoning, and incorporate human feedback via test-time intervention.
+We further propose the token bottleneck models, which allow the users to interpret the evidence selected by Vamos, and speed up inference by nearly 5x.
+Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, and outperforms the best vision-language model by over 66% on EgoSchema.
+Limitations: Although our results show the promise of free-form text-based representations, we believe visual information is still essential for complex video understanding and reasoning.
+We expect future work to investigate alternative visual encoders [66] to extract fine-grained visual information beyond what has been captured by the captions, and to propose pre-training paradigms that better align visual inputs with the input space of LLMs.
+We also believe better benchmarks that require fine-grained, structured visual understanding are needed to rigorously evaluate the impact of representations for video understanding.
 Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology.
-We would like to thank Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their generous feedback on this work.
-Our research was conducted using computational resources at Brown CCV.
+We would like to thank Karttikeya Mangalam, Raiymbek Akshulakov, and Shyamal Buch for their kind help with EgoSchema and ATP; Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their useful feedback and discussions.
+Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.

### v3

@@
 To interpret the important text evidence for question answering, we generalize the concept bottleneck model to work with tokens and nonlinear models, which uses hard attention to select a small subset of tokens from the free-form text as inputs to the LLM reasoner.
-We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, NeXT-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning.
+We evaluate Vamos on five complementary benchmarks, Ego4D, NeXT-QA, IntentQA, Spacewalk-18, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning.
 Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era.
@@
 (b) Overview of the Vamos framework.
-It flexibly unifies distributed visual features and textual video representations such as free-form video captions.
-By performing sequence modeling, Vamos can be applied to multiple video understanding tasks.
+It flexibly unifies distributed visual features and text-based representations such as video captions, and can be applied to diverse video understanding tasks.
 We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [40, 48] on sampled video frames, serve as useful video representations for action modeling from videos?
@@
 TBM naturally supports the incorporation of multimodal information, and allows users to perform causal intervention.
-We perform extensive evaluations on four benchmarks, including the Ego4D dataset [18] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, and EgoSchema [50] for zero-shot long-form video question answering.
+We perform extensive evaluations on five benchmarks, including the Ego4D dataset [17] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, Spacewalk-18 [37] for long-form procedural video understanding, and EgoSchema [50] for zero-shot long-form video question answering.
 We observe that for the direct application of Vamos in the action anticipation task, the text-based representation outperforms its counterpart based on visual embeddings.
-We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 50%.
+We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 66%.
 We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings surprisingly results in marginal or no performance gains.
@@
 Vision-Language Foundation Models.
-Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension.
-Models such as CLIP [59], ALIGN [30], CLOOB [17] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs.
-Another line of vision-language models [43, 10, 45, 63] use a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection.
+Models such as CLIP [59] and ALIGN [29] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss on image and caption pairs.
+Another line of vision-language models [43, 10, 45, 63] combines the masked language modeling objective with image-text contrastive learning, and focus on downstream tasks such as visual question answering, visual commonsense reasoning, and text-guided object detection.
 For videos, VIOLET [16] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling.
@@
 To implement TBM, we design a lightweight token selector as an add-on module for Vamos (Figure 2 middle).
-The token selector takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs.
+It takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs.
 It learns to pick a single token among the candidate tokens $\mathbf{z}_{\text{tvr}}$ by optimizing the objective for a given downstream task.
@@
 For VQA, we evaluate on three datasets:
-EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding.
-Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos.
-It contains 5031 videos and each video has a multiple choice question.
-It requires zero-shot evaluation and releases answers to 500 questions for development.
+EgoSchema [50] is annotated on Ego4D videos for long-form video QA.
+Each video is around 3 minutes and the temporal certificate for humans to solve each task is around 100 seconds.
+It has 5,031 videos and each video is annotated with a multiple-choice question.
+All examples are for zero-shot evaluation.
+Spacewalk-18 [37] is a long-form procedural video understanding benchmark collected on 18 spacewalk videos.
+The total duration is over 96 hours.
+We evaluate on the step recognition task which has a temporal certificate of 140 seconds.
+We follow the zero-shot setup with 1-minute context window, and report step recognition accuracy on the test set.
 NeXT-QA [75] is a popular multiple choice video question answering benchmark that tests video understanding in terms of describing and reasoning the temporal actions.
@@
 We generate zero-shot, task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema.
-We use LLaVA-1.5 to generate captions for NeXT-QA due to its better performance (see Table 4.5 for an ablation study).
+We use LLaVA-1.5 to generate captions for NeXT-QA and Spacewalk-18 due to its better performance (see Table 4.5).
 For Ego4D LTA, we sample the center frame for each video segment to generate its caption.
@@
 For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, and then generate and concatenate the frame-level captions.
-We sample 6, 6, and 12 frames for NeXT-QA, IntentQA, and EgoSchema, respectively.
+We sample 6, 6, 12, and 12 frames for NeXT-QA, IntentQA, Spacewalk-18, and EgoSchema, respectively.
+For Spacewalk-18, we additional use its provided speech narrations.
 Vamos for Temporal Modeling.
@@
 In this way, we condense a long sequence (644 tokens on average for NeXT-QA) into $k$ tokens.
-TBM is jointly optimized with Vamos during training.
+TBM is jointly optimized with Vamos with LLaMA-3 during training.
 For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA-1.5.
@@
 Table 4.5 shows that the TBM leads to an expected performance drop due to discarding over 90% of the input tokens, and increasing $k$ improves the TBM performance.
-When $k=40$, Vamos achieves a competitive 65.3% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
+When $k=40$, Vamos achieves a competitive 69.6% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
 Figure 4: Illustration of the token bottleneck model and the impact of intervention on model’s predictions.
@@
 Input Selected tokens ($k$) Cau. Tem. Des. All
-vision all 69.6 67.2 74.7 69.6
-text all 75.5 71.3 81.1 75.0
-text 20 / 644 59.5 56.8 64.6 59.5
-text 40 / 644 65.4 62.4 70.5 65.3
+vision all 71.9 67.4 75.6 71.0
+text all 77.2 75.3 81.7 77.3
+text 20 / 644 68.1 64.6 70.8 67.4
+text 40 / 644 70.1 67.4 72.2 69.6
 Visualization of Vamos predictions.
@@
 The two positive examples demonstrate Vamos’ ability to select tokens that serve as evidence for question answering.
-The negative example is challenging: even highly related tokens such as “race” and “track” are selected by Vamos, it still fails to reason the purposes of the tyres in the video.
+The negative example is challenging: even highly related tokens such as “snow” and “glide” are selected by Vamos, it still fails to reason the correct activities in the video.
 Test Time Intervention.
@@
 We study the impact of LLMs on the EgoSchema benchmark.
-As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines.
-This again demonstrates that Vamos can directly benefit from advances in the large language models, making its decoupled perception and reasoning design desirable for practitioners.
+As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines, which again demonstrates that Vamos can directly benefit from advances in LLMs.
 4.7 Comparison with State-of-the-art
-Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 4.74.74.7, and 4.7.
+We compare our proposed Vamos with other state-of-the-art models in Tables 4.74.74.7, and 4.7.
 We train Vamos with LLaMA2-7B on Ego4D LTA and LLaMA3-8B on NeXT-QA and IntentQA.
@@
 On Ego4D LTA, Vamos outperforms previous works, without relying on domain-specific video encoders [47].
+Finally, on Spacewalk-18, we use GPT-4o as the LLM and achieves 18.6% accuracy, which significantly outperforms the prior best zero-shot performance of 13.6%.
 Table 7: Comparison on NeXT-QA benchmark. * with additional supervision.
@@
 Vamos (ours) 77.2% 75.3% 81.7% 77.3%
-Table 8: Comparison with SOTA on IntentQA
+Table 8: Comparison with SOTA on IntentQA.
  Model CW CH TP&TN ALL
@@
 We further propose the token bottleneck models, which allow the users to interpret the evidence selected by Vamos, and speed up inference by nearly 5x.
-Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, and outperforms the best vision-language model by over 66% on EgoSchema.
+Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, Spacewalk-18, and outperforms the best vision-language model by over 66% on EgoSchema.
 Limitations: Although our results show the promise of free-form text-based representations, we believe visual information is still essential for complex video understanding and reasoning.
@@
 We also believe better benchmarks that require fine-grained, structured visual understanding are needed to rigorously evaluate the impact of representations for video understanding.
-Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology.
+Acknowledgements
+This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology.
 We would like to thank Karttikeya Mangalam, Raiymbek Akshulakov, and Shyamal Buch for their kind help with EgoSchema and ATP; Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their useful feedback and discussions.
 Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
+Appendix 0.A Additional Experiments and Results
+Textual representation format.
+On the Ego4D LTA task, we observe that explicitly regularizing the text-based inputs to contain only verbs and nouns lead to slightly improved performance (0.878 versus 0.890 action edit distance).
+We hypothesize that this is due to the nature of the task, which focuses solely on predicting future verbs and nouns.
+In Table 0.A, we again compare the action-based and caption-based text representations, but on the EgoSchema zero-shot VQA task.
+In contrast to the LTA performance, the caption-based representation performs much better, intuitively because solving the video question answering would require more details about the video, which can be provided by the general-purpose captions.
+Frame number on EgoSchema.
+In Table 3 we compare different frame numbers and the zero-shot performance on EgoSchema.
+We uniformly sample 1, 4, and 12 frames from the videos and concatenate the captions of each frame to form the text video representations.
+We use GPT-3.5 as the reasoner.
+The results show that using more frames leads to better performance as more information and temporal evidence are provided.
+Temporal ordering information.
+We investigate the influence of temporal information for long-form video understanding by shuffling the frame order when concatenating captions.
+In Table 0.A, we compare the ordered and shuffled 12-frame captions.
+Surprisingly, the performance drop (2%) by shuffling the video captions is not as significant as we expect.
+We suspect that LLMs may have strong capability of “auto-correcting” the order of the input sentences, even if they are shuffled.
+Table A1: Comparison of action and caption for zero-shot VQA on Egoschema.
+Model Input Type Full Set Acc.
+GPT-4 action 38.12%
+GPT-4 caption 48.26%
+Table A2: Ablation on the influence of temporal information. 12 frames are sampled.
+Shuffle   Full Set Acc.
+39.22%
+41.24%
+Appendix 0.B Addition Implementation Details
+Vamos.
+We train Vamos on 4 A6000 GPUs for 2, 5, 10 epochs on Ego4D, IntentQA, and NeXT-QA respectively.
+For NeXT-QA, we set the maximum sequence length of pure vision input as 128, and 1200 for captions and captions + vision inputs, for IntentQA, the sequence lengths are set as 128 and 512 respectively.
+Token bottleneck model.
+For the token bottleneck model (TBM), we use a 2-layer transformer encoder with 2 attention heads, and a hidden size of 128.
+No additional positional encoding is added by the token bottleneck model.
+In order to condense the input sequence for interpretability and select the most relevant tokens, the TBM is task dependent, namely taking the questions and the candidate answers (when available) as inputs along with the video captions.
+The selected tokens, as opposed to their corresponding encoded embeddings by the token bottleneck, are fed into the LLM, as illustrated in Figure 2 (middle) in the main paper.
+Appendix 0.C Additional Visualization
+Figure A3 illustrates the input format for long-term action anticipation and video question answering tasks.
+In the LTA task, the task-specific inputs are the discrete action labels and the target output is a future action sequence.
+For supervised VQA, the task-specific input is composed of instructions, video representations (vision and text), question and choices.
+The target output is the chosen answer to the question.
+For zero-shot VQA on EgoSchema, the video representation only consists of the text descriptions.
+Figure A3 shows the prompt designs for all tasks.
+Figure A4 shows four example video captions generated by LLaVA and BLIP-2 on the NeXT-QA dataset, respectively.
+We observe that in general the LLaVA-generated captions are longer and more detailed.
+Recall that Table 5 shows that the more detailed LLaVA captions also lead to higher VQA performance.
+In Figure A1 and Figure A2, we provide additional prediction examples from Vamos and Vamos with the token bottleneck model respectively.
+Figure A1: More examples of Vamos video question answering and manual intervention.
+Figure A2: Additional illustration of predictions with TBM and manual intervention.
+The token bottleneck selects tokens highly related to the question and intervention corrects wrong predictions without training.
+Figure A3: Input of Vamos for long-term action anticipation and video question answering. <v> in prompts denotes vision tokens.
+Figure A4: Example of video captions generated by BLIP-2 and LLaVA on the NeXT-QA dataset, respectively.